Fermi National Laboratory
 

Operation Summary for January 29, 2010 - February 1, 2010

 

The Integrated Luminosity for the period from 1/25/10 to 2/1/10 was 15.03 pb-1.
NuMI reported receiving 4.2E18 protons on target during this same period.

Friday Jan. 29

The day shift began with the Tevatron (TeV) recovering from a quench, with Linac personnel repairing a LCW leak on a Klystron RF station (KRF2), with the Recycler, Antiproton Source, Booster, Main Injector, NuMI, MiniBooNE, and MTest experiments waiting for the KRF2 repair, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiment waiting for the next store.

During the day shift, the TeV recovery continued. Operations put the TeV through a dry squeeze while D0 personnel made an access into their collision hall. Linac personnel repaired the KRF2 LCW leak; Linac resumed sending beam at 11:30 AM. The D0 low beta magnets quenched during a dry squeeze and a TeV quadrupole power supply (D0Q6) tripped off with a ground fault indication. Humidity levels increased in the vicinity of D0Q6 and the TeV water loop indicated an LCW leak. TeV personnel accessed and found the bolted joint for the D0Q6 warm bus connection had been compromised. Both ceramic connections were broken, spraying LCW onto the magnet and a local ion pump gauge. Mechanical support personnel began repairs. A MI-10 quadrupole power supply failed. The MI ramp was halted in order for EE Support personnel to investigate. They replaced an SCR gate firing fanout board and MI resumed sending beam by 4 PM.

During the evening shift, mechanical personnel completed their repair of D0Q6 at 6:30 PM. TeV turn on and checkout followed. Operations put the TeV into a HEP shot setup at 9:30 PM. The shot setup ended with a quench of the D0 low betas due to a D0Q2 lead failure. Operations called in EE Support personnel to investigate. Beam was delivered to the Pbar Source, NuMI, MiniBooNE, and MTest when available. Pbar transferred antiprotons to the Recycler when appropriate.

Saturday Jan. 30

The midnight shift began with TeV and cryo system personnel investigating the quench, with Operations monitoring the stash (59.6E10), with Pbar stacking (19.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, TeV personnel continued to investigate the lead failure from the previous shift. EE Support personnel came in to troubleshoot upstairs. No obvious problems were found; TeV personnel accessed to investigate in the tunnel. Technicians reconnected and reterminated the lead voltage connections and the TeV turn on followed. Just prior to ramping, TeV sector B3 quenched due to bad inductance settings for the B32 cell. Operations recovered the TeV and successfully completed a dry squeeze and two rounds of orbit checks.

During the day shift, Operations had the TeV in shot setup by 8:45 AM, but it came to an abrupt halt when a TeV kicker (AAK1) pre-fired causing a messy abort that quenched several houses. Recovery lasted several hours. Unfortunately, the TeV magnets at BA and BB quenched unexpectedly during the TeV checkout and turn on. The TeV A11 Kautzky valve developed an electrical problem and would not close. As a result, cryo system techs accessed the Transfer Hall to make repairs. The TeV abort kickers were put in conditioning mode during the access.

During the evening shift, cryo techs had the Kautzky valve at A11 repaired by 9 AM; TeV recovery and turn on followed. A flaky water temperature interlock for the Gradient Magnet Power Supply (GMPS) prompted a Booster access to investigate.

Sunday Jan. 31

The midnight shift began with a Booster access, with Operations monitoring the stash and stack, and with NuMI and MiniBooNE waiting on Booster beam.

During the midnight shift, the TeV A2 dump would not make up following a dump test. EE Support personnel responded and replaced a failed breaker. Operations successfully conducted a TeV QBS ramp at 1:30 AM. Booster personnel jumpered the errant GMPS interlock and beam operations resumed by 3 AM. Operations checked the TeV orbits and put the TeV into shot setup at 3:45 AM. The TeV B11 vertical separator sparked during the squeeze and ended another shot setup prematurely. Operations put the separator into conditioning mode while the cryo systems recovered; Operations checked out the TeV and prepared for another turn on.

During the day shift, Operations completed the TeV checkout, turn-on, and orbit checks and had the TeV in shot setup by 12:15 PM. Operations established store 7559 at 1:30 PM with an Average Initial Luminosity (AIL) of 227E30.

During the evening shift, at 5 PM, the B11 vertical separator sparked again and caused E1 to quench and then a Kautzky valve at E1 stuck in its open position; cryo techs accessed the TeV for repairs. Operations put the B11 vertical separator into conditioning mode for the rest of the shift. Cryo recovery was well underway by the end of the shift.

Monday Feb. 1

The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TeV checkout and turn on continued from the previous shift. The TeV E1 cryo system recovery proceeded slowly due to issues with the vacuum. To make matters worse, once the E1 cryo system was ready to initiate the cold compressor pumpdown, the compressor would not start up and needed to be replaced. Operations called in cryo techs to make the necessary repairs. The cryo techs had the E1 cryo system recovered by 2:15 AM. Operations had the TeV ramping by 2:30 AM. Operators took the B11 separator out of conditioning mode and checked the orbits prior to shot setup, which began at 4:15 AM. Operations established store 7562 at 5:45 AM with an AIL of 259E30.

The Plans for Monday

The Run Coordinator will allow Pbar personnel to access their water cage and for Recycler personnel to conduct studies. The plans for the rest of the day are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory